# BubbleONet: A Physics-Informed Neural Operator for High-Frequency Bubble Dynamics

### v1

@@ 5.2.2 Cross validation for generalization
Some validation cases, such as the bubble simulated with a frequency of $900~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$, exhibit noticeable discrepancies between the ground truth and the model prediction, as shown in Figure. 8. This may be due to the lack of similar features in the training data, leading to poor generalization in this particular case. To address this, a 5-fold cross-validation (CV) was conducted to enhance the model’s robustness.
A comparison between the results of the directly trained model and the model trained with CV reveals a significant improvement. Specifically, the absolute error range decreases by an order of magnitude, from [0.6,0.6][-0.6,0.6] to approximately [0.06,0.06][-0.06,0.06]. Notably, the frequency domain results indicate that even the directly trained model exhibits sufficient phase and frequency resolution to successfully capture the driving frequency, demonstrating the effectiveness of the rowdy activation function. However, the model fails to accurately predict the magnitude of the driving frequency component,i.e., it does not predict how strong the frequency response is, resulting in a mismatch between the predicted and true amplitudes. These findings suggest that the primary source of the discrepancy is likely due to data sparsity, rather than a fundamental inability to learn the frequency characteristics. As such, applying cross-validation helps alleviate this issue by improving the model’s generalization performance in underrepresented regions of the parameter space.
@@
Figure 8: Effect of cross-validation on model generalization. The case of a bubble driven at $900~\text{kHz}$ and $9\times 10^{5}~\text{Pa}$ shows significant improvement in prediction accuracy after applying 5-fold cross-validation.
5.2.3 Interpolation and extrapolation
To further investigate the model’s generalization capability, both interpolation and extrapolation tests are conducted. As shown in Table 3, the data for this analysis are simulated using amplitudes of $5.5\times 10^{5}~\text{Pa}$ and $11\times 10^{5}~\text{Pa}$, where the former lies within the range of the training data (interpolation), and the latter is outside the training range (extrapolation). The corresponding driving frequencies span from $600~\text{kHz}$ to $2500~\text{kHz}$, with intervals of $100~\text{kHz}$, resulting in a total of 40 datasets for this study.
Table 3: Design of experiments for interpolation and extrapolation
Parameters Range Numbers of Sampling
Pressure amplitude, amp (Pa) $5.5,11]\times 10^{5}$ 2
Pressure frequency, $f$ (kHz) $[600,2500]$ 20
Figure 9: Time-domain and frequency-domain of interpolation results for bubbles simulated at $600~\text{kHz}$, $1000~\text{kHz}$, $1300~\text{kHz}$, and $2000~\text{kHz}$, with a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$.
The model used in this analysis is the optimized version obtained after applying cross-validation. Interpolation results at an amplitude of $5.5\times 10^{5}~\text{Pa}$ are presented in Figure. 9, with four representative frequencies: $600~\text{kHz}$, $1000~\text{kHz}$, $1300~\text{kHz}$, and $2000~\text{kHz}$.
  • 1.
    For the case of $600~\text{kHz}$, the model performs well during the first half of the time series, but noticeable errors appear in the second half. The predicted radius exhibits reduced oscillation compared to the ground truth, which is also reflected in the frequency domain where the magnitude at the driving frequency is underestimated. This underperformance is generally observed in the range of $600~\text{kHz}$ to $800~\text{kHz}$.
  • 2.
    When the frequency increases to the range of $900~\text{kHz}$ to $1200~\text{kHz}$, the model predictions align well with the ground truth in both the time and frequency domains. For example, the case at $1000~\text{kHz}$ achieves an absolute error within the range of [0.02,0.02][-0.02,0.02].
  • 3.
    Prediction quality degrades again for frequencies between $1300~\text{kHz}$ and $1500~\text{kHz}$. The case at $1300~\text{kHz}$ shows clear discrepancies in both the time domain and the spectral magnitude at the driving frequency.
  • 4.
    However, when the frequency increases further to the range of $1600~\text{kHz}$ to $2000~\text{kHz}$, the model’s performance improves significantly. For instance, the case at $2000~\text{kHz}$ demonstrates excellent agreement with the ground truth, with an absolute error confined within (0.01,0.01)(-0.01,0.01).
Such periodical learning feature is likely due to sinusoidal function introduced in Rowdy activation function.
@@ 5.2.3 Interpolation and extrapolation
Figure 10: Time-domain and frequency-domain of extrapolation results for bubbles driven at $2100~\text{kHz}$ and $2500~\text{kHz}$, with a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$.
@@
Two extrapolation cases at an amplitude of $5.5\times 10^{5}~\text{Pa}$ are presented in Figure. 10, corresponding to driving frequencies of $2100~\text{kHz}$ and $2500~\text{kHz}$. Since these cases fall outside the function space used for operator learning, it is anticipated that the model may struggle to generalize. Indeed, both cases fail to accurately capture the driving frequency in the frequency domain. Nevertheless, a recurring learning pattern remains observable. In particular, the case at $2500~\text{kHz}$ exhibits smaller errors in the time domain and reduced discrepancies in the frequency spectrum. The corresponding MSE for each case are shown in Figure. 11 with blue color representing the interpolation and orange color representing the extrapolation, where a similar periodic pattern can be observed, further illustrating the model’s frequency-dependent behavior.
Figure 11: Mean square error across cases with frequencies ranging from $600~\text{kHz}$ to $2500~\text{kHz}$, at a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$. Blue bars represent the frequencies learned by the model, while orange bars indicate frequencies outside the training domain.
Figure 12: Time-domain and frequency-domain extrapolation results for bubbles driven at $1000~\text{kHz}$ and $2000~\text{kHz}$, with a fixed amplitude of $11\times 10^{5}~\text{Pa}$.
For the extrapolation in terms of amplitude, two cases with an amplitude of $11\times 10^{5}~\text{Pa}$ at frequencies of $1000~\text{kHz}$ and $2000~\text{kHz}$ are shown in Figure. 12. In both cases, the absolute errors are reasonable ranging approximately within [1,1][-1,1] and [0.4,0.4][-0.4,0.4], with corresponding MSE of $0.56\times 10^{-3}$ and $4.10\times 10^{-3}$, respectively. However, the driving frequencies are still accurately captured in the frequency domain. This suggests that the source of error is primarily due to the amplitude, as the model has not encountered similar features during training. These findings further confirm the sensitivity of the model to unseen amplitude ranges.
Another extrapolation study is conducted for a longer simulated time range, extending the original duration from $55~\mu\text{s}$ to $100~\mu\text{s}$. The results are shown in Figure. 13. In this scenario, the model fails to make accurate predictions, even within the time interval it was previously trained on, i.e., from 0 to $55~\mu\text{s}$. Beyond this point, the predicted radius begins to shrink significantly, indicating that the model has entered a region outside of its learned function space. Further evidence of this failure is seen in the frequency domain, where the model is unable to capture either the natural frequency or the driving frequency, let alone their magnitudes. These results highlight the model’s limitation in generalizing to extended temporal domains.
@@
Figure 13: Time-domain and frequency-domain extrapolation results for bubble dynamics over $100\mu s$ driven at $600~\text{kHz}$ and $2000~\text{kHz}$, with a fixed amplitude of $1\times 10^{5}~\text{Pa}$.
5.3 Two-step training for K-M eqaution with single initial radius
The core idea of DeepONet is to construct the “basis” functions using the trunk network, while the branch network learns the corresponding “coefficients.” To enhance the performance of DeepONet, Lee et al. [56] proposed a novel two-step training strategy. In their approach, QR factorization is applied to the pretrained “basis” generated by the trunk network, followed by training the branch network. Building upon this framework, Peyvan et al. [14] further improved the method by employing singular value decomposition (SVD) as a replacement for QR decomposition. Following this idea, we applied the two-step training technique to BubbleONet. For more details on these applications, readers are referred to [56, 14].
Figure 14: Two step training architecture for BubbleONet.
The training process and network architecture are illustrated in Figure. 14. In the first step, a learnable variable $A_{m}$ is introduced to mimic the latent representation from the branch network in single-step training. The trunk network is then trained to determine its optimal hyperparameters $\phi^{*}$. Subsequently, the branch network is trained using the SVD-transformed output of the pretrained trunk network to find the optimal hyperparameters $\theta^{*}$ for the branch network. Both the trunk and branch networks share the same architecture: each consists of 8 layers with 512 neurons per layer. Rowdy activation function is applied to both branch net and trunk net with two sinusoidal functions, initialized identically as in Eq. 7. The data, optimizer and other hyper parameters are the same as the case in section 5.2. The convergence history for both the trunk and branch networks is shown in Figure. 15. Meanwhile, a comparison of training time using different training technique shown in Table. 4. It can be obeserved that the two-step training is $20\%$ faster compared to the single-step training process.
Figure 15: Convergence history of trunk net and branch net in two-step training.
Table 4: Comparison of training time for single-step and two-step training
Training Time(h)
Single-step 72
Two-step 57.5
Validation results are presented in Figure. 16, which includes three representative cases. Cases (a) and (c) share the same parameters as Figure. 7(a) and Figure. 7(c), respectively: (a) corresponds to a low-frequency case with a frequency of $523~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$, while (c) represents a high-frequency case with a frequency of $1927~\text{kHz}$ and an amplitude of $8\times 10^{5}~\text{Pa}$. Both cases are well predicted, with mean square error of $4.93\times 10^{-3}$ and $2.79\times 10^{-4}$, respectively. Figure. 16(b) corresponds to the same case as in Figure. 8, with a frequency of $900~\text{kHz}$ and a pressure amplitude of $9\times 10^{5}~\text{Pa}$. The bubble profile is accurately predicted even without cross-validation.
Figure 16: Validation of two-step training method for bubbles simulated using K-M equation with: (a) a frequency of $523~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$, (b) a frequency of $900~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$, and (c) a frequency of $1927~\text{kHz}$ and an amplitude of $8\times 10^{5}~\text{Pa}$.
Interpolation is then conducted to evaluate the inference capability of the model. Figure. 17 shows four different frequencies: $600~\text{kHz}$, $1000~\text{kHz}$, $1300~\text{kHz}$, and $2000~\text{kHz}$ for a pressure magnitude of $5.5\times 10^{5}~\text{Pa}$. The corresponding MSEs are presented in Figure. 18. Similar periodical learning feature is observed but is different from the results in the previous section, where the number of sinusoidal functions is reduced in two-step training technique. Performance at lower frequencies, such as $600~\text{kHz}$ and $700~\text{kHz}$, is worse compared to the single-step training, whereas performance at other frequencies improves, with reduced error, though still exhibiting periodic learning behavior. On the other hand, extrapolation to unseen frequencies results in higher errors than single-step training, which may be attributed to the reduction of sinusoidal components in the rowdy activation function. An example of extrapolation at $2400~\text{kHz}$, corresponding to the minimum MSE in that range, is shown in Figure. 19.
Figure 17: Time-domain and frequency-domain of interpolation results from two-step trained model for bubbles simulated at $600~\text{kHz}$, $1000~\text{kHz}$, $1300~\text{kHz}$, and $2000~\text{kHz}$, with a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$.
Figure 18: Mean square error across inference cases from two-step trained model with frequencies ranging from $600~\text{kHz}$ to $2500~\text{kHz}$, at a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$. Blue bars represent the frequencies learned by the model, while orange bars indicate frequencies outside the training domain.
Figure 19: Time-domain and frequency-domain of extrapolation results of two-step trained model for bubbles driven at $2400~\text{kHz}$, with a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$.
[…]
The final set of extrapolation tests focuses on the bubble excitation time, in comparison with Fig. 13 as shown in Fig. 21. Using a pressure profile with an amplitude of $1\times 10^{5}~\text{Pa}$ and frequencies of $600~\text{kHz}$ and $2000~\text{kHz}$, the simulation time is extended from $55~\mu s$ to $100~\mu s$. Although the two-step trained model performs better than the single-step model, it still fails to accurately predict the bubble dynamics at unseen time durations.
Figure 20: Time-domain and frequency-domain extrapolation results from two-step trained model for bubbles driven at $1000~\text{kHz}$ and $2000~\text{kHz}$, with a fixed amplitude of $11\times 10^{5}~\text{Pa}$.
Figure 21: Time-domain and frequency-domain extrapolation results from two-step trained model for bubble dynamics over $100\mu s$ driven at $600~\text{kHz}$ and $2000~\text{kHz}$, with a fixed amplitude of $1\times 10^{5}~\text{Pa}$.
@@ 5.3 Two-step training for K-M eqaution with single initial radius
5.4 Two-step training for K-M equation with multi initial radii
@@
In this section, we focus on the high-dimensional basis constructed by the trunk network, as illustrated in Figure.  22. Given $n$ time sampling points and $k$ different initial radii $R_{0}$, meshgrids of size $n\times k$ (time and radius) are generated for each $R_{0}$, resulting in $k$ meshgrids. These are flattened and stacked into a single input of size $(n\times k)\times 2$ for the trunk network. The branch network receives pressure profiles of size $m\times n_{p}$, where $m$ is the batch size and $n_{p}$ is the number of pressure sampling points. The model outputs predictions of size $m\times(n\times k)$, which are reshaped to $m\times k\times n$, representing bubble dynamics over time for each initial radius.
Figure 22: Implementation of multi-initial radii scenario, with initial radius and time as the input to trunk net.
@@
To showcase the capability, five initial radii, $50\,\mu m$$60\,\mu m$$70\,\mu m$$80\,\mu m$ and $90\,\mu m$, are considered. Pressure amplitudes and frequencies are sampled within the same ranges specified in Table 1, with 10 amplitude values and 30 frequency values, resulting in 300 distinct pressure profiles. For each initial radius, 300 bubble profiles are generated, totaling 1500 samples. Among these, 80% are used for training and 20% for validation. Consequently, the network architecture is adjusted to account for varying natural frequencies introduced by different initial radii. The Rowdy activation function incorporates four sinusoidal components to better capture these dynamics. The weight for loss function is then changed to $w_{\text{data}=1}$$w_{\text{data}=1000}$, an additional term of initial radii is enforced as the initial condition(IC), with the weight $w_{\text{IC}}=1$
A prediction for bubbles subjected to a pressure oscillation with a frequency of $467~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$ is presented in Figure. 23. In Figure. 23(a), the second peaks of all bubble profiles are highlighted with green circles and magnified on the right for clearer visualization. As shown in Figure. 23(b), all predicted profiles exhibit excellent agreement with the ground truth, with maximum absolute errors of approximately $0.2\,\mu\text{m}$, corresponding to 0.40% for the $50\,\mu\text{m}$ radius and 0.22% for the $90\,\mu\text{m}$ radius.
A second prediction, shown in Figure. 24, considers bubbles subjected to a higher-frequency pressure oscillation of $1800~\text{kHz}$ with an amplitude of $3\times 10^{5}~\text{Pa}$. As highlighted in Figure. 24(a), discrepancies in amplitude between the predicted and true profiles are observed at the second peaks (highlighted by green circles and magnified views). Nevertheless, the maximum absolute errors remain below $0.1\,\mu\text{m}$, corresponding to 0.20% and 0.11% for the $50\,\mu\text{m}$ and $90\,\mu\text{m}$ radii, well within an acceptable range. Furthermore, FFT in Figure. 25 confirms that the model accurately captures both the natural and driving frequencies.
@@
Figure 23: Prediction of bubble dynamics with five different initial radii, under pressure with a frequency of $467~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$. (a) Bubble profile with the second peak highlighted in green circles and shown on the right. (b) Absolute error.

### v3

@@ 5.2.1 Cross validation for generalization
Figure 9: Effect of cross-validation on model generalization. The case of a bubble driven at $954~\text{KHz}$ and $9\times 10^{5}~\text{Pa}$ shows significant improvement in prediction accuracy after applying 5-fold cross-validation.
Some validation cases, such as the bubble simulated with a frequency of $954~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$, exhibit noticeable discrepancies between the ground truth and the model prediction (Figure 9 (left)). This may result from a lack of similar features in the training data, leading to poor generalization in this specific case. To address this limitation, a 5-fold cross-validation (CV) was performed to enhance the model’s robustness.
Comparison between the directly trained model and the CV-trained model shows a marked improvement: the absolute error range decreases by an order of magnitude, from $[0,0.6]$ to approximately $[0,0.06]$. Notably, the frequency-domain results indicate that even the directly trained model provides sufficient phase and frequency resolution to capture the driving frequency, demonstrating the effectiveness of the Rowdy activation function. However, it fails to accurately predict the magnitude of the driving-frequency component-that is, the strength of the frequency response-resulting in a mismatch between the predicted and true amplitudes. These findings suggest that the primary source of discrepancy arises from data sparsity rather than an inherent limitation in learning frequency characteristics. Accordingly, applying cross-validation mitigates this issue by improving the model’s generalization performance in underrepresented regions of the parameter space; see, Figure 9 (right).
5.2.2 Out-of-distribution Results
We perform a systematic study of frequency-domain, time-domain, and pressure-amplitude extrapolation results.
@@ 5.2.2 Out-of-distribution Results
Figure 10: First two columns: Frequency-domain extrapolation results for bubbles driven at $2100~\text{KHz}$ and $2500~\text{KHz}$, with a fixed amplitude of $5.5\times 10^{5}~\text{Pa}$. Right figure: L2-norm errors for different frequencies. Orange bars represents out-of-distribution frequencies.
@@
Two out-of-distribution cases at an amplitude of $5.5\times 10^{5}~\text{Pa}$ are presented in Figure 10 (first two columns), corresponding to driving frequencies of $2100~\text{kHz}$ and $2500~\text{kHz}$. Since these cases fall outside the function space used for operator learning, it is anticipated that the model may struggle to generalize. Indeed, both cases fail to accurately capture the driving frequency in the frequency domain. Figure 10 (Right) shows L2-norm errors for different frequencies.
@@
Figure 11: First and Second Columns: Pressure amplitude extrapolation results for bubbles driven at $1000~\text{KHz}$ and $2000~\text{KHz}$, with a fixed amplitude of $11\times 10^{5}~\text{Pa}$. Third and Forth Columns: Time-domain extrapolation results for bubble dynamics over $100\mu s$ driven at $600~\text{KHz}$ and $2000~\text{KHz}$, with a fixed amplitude of $1\times 10^{5}~\text{Pa}$.
For the amplitude-based out-of-distribution test, two cases with an amplitude of $1.1\times 10^{6}~\text{Pa}$ at frequencies of $1000~\text{kHz}$ and $2000~\text{kHz}$ are shown in Figure 11 (the first and second columns). In both cases, the absolute errors are reasonable, ranging approximately within [0, 0.8][0,\,0.8] and [0, 0.4][0,\,0.4]. However, the driving frequencies are still accurately captured in the frequency domain. This suggests that the source of error is primarily due to the amplitude, as the model has not encountered similar features during training. These findings further confirm the sensitivity of the model to unseen amplitude ranges.
A final out-of-distribution study is conducted for an extended simulated time range, increasing the original duration from $55~\mu\text{s}$ to $100~\mu\text{s}$. The results are shown in Figure 11 (the third and fourth columns). In this scenario, the model fails to make accurate predictions even within the time interval it was previously trained on (i.e., from 0 to $55~\mu\text{s}$). Beyond this point, the predicted radius begins to shrink significantly, indicating that the model has entered a region outside its learned function space. Further evidence of this failure is observed in the frequency domain, where the model is unable to capture either the natural or driving frequency, let alone their magnitudes. These results highlight the model’s limitation in generalizing to extended temporal domains.
@@
5.3 Two-step DeepONet Training for K-M Equation with Multi Initial Radii
@@
In this section, we focus on the two-dimensional basis constructed by the trunk network, as illustrated in Figure 12. Given $n$ time sampling points and $k$ different initial radii $R_{0}$, meshgrids of size $n\times k$ (time and radius) are generated for each $R_{0}$, resulting in $k$ meshgrids. These meshgrids are flattened and stacked into a single input of size $(n\times k)\times 2$ for the trunk network. The branch network receives pressure profiles of size $m\times n_{p}$, where $m$ is the batch size and $n_{p}$ is the number of pressure sampling points. The model outputs predictions of size $m\times(n\times k)$, which are then reshaped to $m\times k\times n$, representing bubble dynamics over time for each initial radius.
Figure 12: Implementation of multi-initial radii scenario.
@@
To demonstrate the model’s capability, five initial radii, $50\,\mu\text{m}$, $60\,\mu\text{m}$, $70\,\mu\text{m}$, $80\,\mu\text{m}$, and $90\,\mu\text{m}$, are considered. Pressure amplitudes and frequencies are sampled within the same ranges specified in Table 1, using 10 amplitude values and 30 frequency values, resulting in 300 distinct pressure profiles. For each initial radius, 300 bubble profiles are generated, yielding a total of 1500 samples. Among these, 80% are used for training and 20% for validation. Consequently, the network architecture is adjusted to account for the varying natural frequencies associated with different initial radii. The Rowdy activation function incorporates four sinusoidal components to better capture these dynamics. The loss function weights are set to $w_{\text{data}}=1$ and $w_{\text{phys}}=1000$, and an additional term enforcing the initial radius as an initial condition (IC) is introduced with a weight of $w_{\text{IC}}=1$.
@@
Figure 13: Prediction of bubble dynamics with five different initial radii, under pressure with a frequency of $467~\text{KHz}$ and an amplitude of $4\times 10^{5}~\text{Pa}$. (a) Bubble profile with the second peak highlighted in green circles and shown on the right. (b) Absolute error.
A prediction for bubbles subjected to a pressure oscillation with a frequency of $467~\text{kHz}$ and an amplitude of $4\times 10^{5}~\text{Pa}$ is presented in Figure 13. In Figure 13(a), the second peaks of all bubble profiles are highlighted with green circles and magnified on the right for clearer visualization. As shown in Figure 13(b), the predicted profiles exhibit excellent agreement with the ground truth, with maximum absolute deviations of approximately $0.3\,\mu\text{m}$ (0.40%) for the $50\,\mu\text{m}$ radius and $0.2\,\mu\text{m}$ (0.22%) for the $90\,\mu\text{m}$ radius.
@@
A second prediction, shown in Figure 14, considers bubbles subjected to a higher-frequency pressure oscillation of $1800~\text{KHz}$ with an amplitude of $3\times 10^{5}~\text{Pa}$. As highlighted in Figure 14(a), discrepancies in amplitude between the predicted and true profiles are observed at the second peaks (highlighted by green circles and magnified views). Nevertheless, the maximum absolute errors remain below $0.061\,\mu\text{m}$, corresponding to 0.12% and 0.07% for the $50\,\mu\text{m}$ and $90\,\mu\text{m}$ radii, well within an acceptable range. Furthermore, FFT in Figure 15 confirms that the model accurately captures both the natural and driving frequencies.
5.3.1 Ablation Study
Table 3: Ablation study of Rowdy terms in the Two-Step DeepONet under a pressure amplitude of $1\times 10^{6}\,\text{Pa}$ and a frequency of $2000\,\text{kHz}$.
Model L2-norm Error
2-Step DeepONet 0.287%
2-Step DeepONet + 1 term (Rowdy) 0.043%
2-Step DeepONet + 2 terms (Rowdy) 0.020%
2-Step DeepONet + 3 terms (Rowdy) 0.051%
An ablation study is conducted to assess the effectiveness of the Rowdy activation terms; see Table 3. Four experiments are performed using the two-step DeepONet architecture with varying numbers of Rowdy activations, validated against bubble dynamics subjected to a pressure amplitude of $1\times 10^{6}\,\text{Pa}$ and a high driving frequency of 2000 kHz. When no Rowdy activation is employed, the model captures only the natural frequency of the bubble but fails to reproduce the driving frequency, as also reported in Ref. [zhang2025bubble], resulting in the highest prediction error of 0.287%. As the number of Rowdy terms increases, the model’s accuracy improves and the prediction error decreases. However, when three Rowdy terms are introduced, overfitting occurs, leading to a slight increase in error.
5.4 Two-step DeepOKAN with Multi Initial Radii
Spectral bias was introduced by Rahaman et al. [rahaman2019spectral] in the context of deep networks where these networks in spite of following the universal approximation theorem solely aims to approximate the low frequency components of a function often neglecting the essential high frequency features present in an underlying function. The previous section dealt with overcoming the spectral bias while learning the high frequency bubble dynamics using two-step DeepONet framework with the help of Rowdy activation functions. Likewise, several recent developments have focused on addressing the spectral bias mainly within the context of MLP; due to the nascency of KAN there has not been any work that tackled spectral bias using the KAN architecture. The two-step DeepOKAN model was trained on an NVIDIA V100 GPU with 64 GB of system RAM.
Figure 16: The training performance of different basis functions for trunk network training for high frequency bubble dynamics. Here, NN in the context of radial basis functions (RBF) means the number of functions at a given node. Spline basis has polynomial order ($k$) = 2 and grid size ($G$) = 40. The training undergoes early stopping under convergence.
Figure 17: The performance of different basis functions while learning the bubble dynamics under pressure with a frequency of $491$ kHz and amplitude of $8\times 10^{5}$ Pa. (left) RBF with $N=6$ (middle) RBF with $N=25$ (right) Spline with $k=2$ and $G=40$. Note that these are intermediate results ($\tilde{R}(\phi)$) solely from trunk net, $\tilde{R}(\phi)=A_{m}T(\phi)^{T}$ to justify the utilization of a specific formulation of the basis function within the KAN architecture.
Figure 16 showcases the training performance of the trunk network for different basis functions utilized in the KAN framework. Here, we note that spline basis with continual learning outperforms all the other configurations of the basis functions. Increasing the number of funtions $N$ from 6 to 25 for the RBF [abueidda2025deepokan] basis is observed to have negligible effects on the training performance with added computational cost incurred from over-parametrization. In a way, this is confirmed by the results shown in Figure 17 where RBF basis unlike the spine basis fails to learn the internal high frequency features of the bubble dynamics even at lower frequency ranges. Although, Figure 16 suggests almost similar training performance for spline basis with and without continual training, it is does not reveal the actual improvement. Figure 18 showcases this improvement for a high frequency, high pressure scenario where the spline basis without continual learning only captures the localized global frequencies while avoiding the high-frequency features of the bubble dynamics. On the other hand, with the use of continual training, we observe that the model captures all the dominant frequencies, with an order-of-magnitude improvement in the training error, as shown in Figure 18.
Figure 18: Comparison of the training performance of trunk network based on spline basis with continual learning and without continual learning while learning bubble dynamics under pressure with a frequency of $1928$ kHz and amplitude of $11\times 10^{5}$ Pa. Note that these are intermediate results ($\tilde{R}(\phi)$) for trunk net, $\tilde{R}(\phi)=A_{m}T(\phi)^{T}$.
Based on the above observations, it can be inferred that KAN as an independent architecture like MLP, may not be able to overcome spectral bias solely by itself. However, with a slight modification in the training via continual learning we are able to overcome the spectral bias without the utilization of any additional trainable parameters via activation functions unlike the two-step DeepONet discussed in the previous section. We also note that this is made possible through the internal parametrization of the basis functions used in the KAN architecture that substantially improves the representation capacity of the KAN-based framework in the high frequency setting. However, an equivalent MLP-based model do not have this favorable characteristic and requires additional parameterizations through adaptive or Rowdy activation functions to overcome the spectral bias as shown in the previous sections. It is also worthwhile to note that, in scenarios where the segregation of the dataset based on their dominant frequency is not feasible prior to model training, it may not be possible to implement continual learning. In such cases one may have to devise alternate strategies to overcome spectral bias using KAN or resort to using MLP with adaptive activation functions to overcome the spectral bias as discussed in the previous sections.
5.4.1 Inference
Table 4: Hyperparameters for 2-step DeepOKAN
Network Learning rate Basis Architecture
Trunk 0.0001 Spline ($k=2$, $G=40$) [2, 150, 150, 150, 150]
Branch 0.0001 RBF ($N=6$) [2000, 150, 150, 150, 150, 150, 150]
While the trunk network utilizes spline basis over the RBF [abueidda2025deepokan] basis for reasons justified in the previous subsection, we utilize the RBF basis for the branch network mainly owing to its computational efficiency, leading to reduced training time compared to the spline basis. The architecture for trunk and branch networks is shown in Table 4. Figure 19 and Figure 20 provide an overview of the two-step DeepOKAN model during inference. For a fixed amplitude of pressure, it can be observed that the model captures the localized dominant high frequency features along with the low frequency bubble dynamics for varying magnitudes of frequency. However, for extremely large frequencies such as $1800$ kHz, it can be observed that although the model predicts the dominant frequency with good accuracy, it poorly approximates the higher frequencies as shown in Figure 20. This may be improved further through continual learning of the trunk network where such extremely high frequencies may be treated as a separate batch (180020001800-2000 kHz) which can be learned in progression after the model learns a specific batch of frequencies with comparatively lower magnitudes (<1800<1800 kHz).
Figure 19: Validation of two-step training method for bubbles simulated using K-M equation with: a frequency of (a) $527$ kHz, (b) $873$ kHz, and (c) $1473$ kHz for a fixed amplitude of $1\times 10^{6}$ Pa.
Figure 20: Validation of two-step training method for bubbles simulated using K-M equation with: a frequency of $1800$ kHz and an amplitude of (a) $4\times 10^{5}$ Pa, (b) $6\times 10^{5}$ Pa, and (c) $1\times 10^{6}$ Pa.
5.4.2 Comparison with vanilla DeepOKAN
Here, we compare the results for the multi-radii case with the vanilla DeepOKAN [abueidda2025deepokan] framework. While Liu et al. [liu2024kan] proposed B-splines as the potential basis for the KAN architecture, Abueidda et al. [abueidda2025deepokan] proposed DeepOKAN with the RBF basis owing to their superior approximation properties and computational efficiency. Here, we compare the multi-radii two-step DeepOKAN with the vanilla DeepOKAN complying with the layout shown in Figure 12 and the architecture shown in Table 4 (with RBF basis) for fair comparison of both the architectures.
Figure 21: Comparison of (left) DeepOKAN [abueidda2025deepokan] and (right) 2-step DeepOKAN for bubble dynamics simulated using K-M equation at a frequency of $1800$ kHz and an amplitude of $1\times 10^{6}$ Pa.
Figure 21 shows the comparison of both the architectures for a high frequency, high pressure test case. Here, we note a clear degradation of performance using the DeepOKAN [abueidda2025deepokan] framework in the multi-radii setting and the utilization of the RBF basis doesn’t favor the operator learning framework in the context of high frequency bubble dynamics. This observation also complies with the results in Figure 17 where it was shown that the performance of spline basis surpasses the RBF for the trunk network in the two-step DeepOKAN framework. To this end, we emphasize the need for careful selection of a suitable basis function based on the type of problem being solved as opposed to selecting a universal basis for the KAN architecture. Here, we also note the benefit of using a two-step operator learning framework which allows independently tuning the trunk and branch networks after taking into account factors like accuracy and computational efficiency of different basis functions as opposed to training them together as a single framework at least in the context of KAN. Furthermore, while the single step operator learning frameworks have showcased exceptional performance for various types of problems, it may be beneficial to transition to a two step operator learning framework based on the level of difficulty (such as overcoming the spectral bias shown in the current work). There are also exceptions to this. For example, FNO [li2020fourier] operates directly in the frequency domain and can effectively address issues related to spectral bias in that domain; however, it is difficult to implement for high-dimensional problems. The comparison of the proposed two-step operator frameworks with such state-of-the-art neural operators will be compared and discussed in the next section.
5.5 Comparison with Different Neural Operators
A fair benchmark study is conducted to compare several operator learning models, including Two-Step DeepONet, the Fourier Neural Operator (FNO) [li2020fourier], the Wavelet Neural Operator (WNO) [tripura2023wavelet], the Operator Transformer (OFormer) [li2022transformer], the Convolutional Neural Operator (CNO) [raonic2023convolutional] and the proposed Two-Step DeepOKAN. Training is performed using bubble dynamics data corresponding to a single initial radius simulated by the Keller–Miksis equations. Validation is carried out across three different pressure amplitudes, $1\times 10^{5}$, $6\times 10^{5}$, and $1\times 10^{6}\,\text{Pa}$, and three driving frequencies of 600, 1300, and 2000 kHz, representing low, medium, and high parameter regimes, respectively.
Table 5: Comparison of relative $L_{2}$ errors (mean $\pm$ standard deviation over 10 realizations) for different neural operator models of bubble dynamics under varying frequencies and pressure amplitudes. The best results are highlighted in bold.
Amp (Pa) Freq (KHz) FNO WNO OFormer CNO 2-Step DeepONet 2-step DeepOKAN
[…]
As shown in Table 5, all six models are evaluated over ten independent initializations, and their ensemble performance is compared using the distribution of $L_{2}$-norm errors (mean $\pm$ standard deviation). The two-step DeepOKAN demonstrates the highest accuracy and precision in most cases, followed by the two-step DeepONet. The results indicate that the two-step DeepOKAN maintains consistent performance in most cases within the specified frequency and pressure ranges, as summarized in Table 5.